OpenStudy (anonymous):

The area of a triangle is found using the following formula:

OpenStudy (juanpablojr):

Options?

OpenStudy (anonymous):

There are 2 formula's for that: If the base and height is known then: \[A = \frac{ 1 }{2 } \times base \times height\] And if the length of the 3 sides of the triangles is known than: Let side1 be a side2 be b and side 3 be c \[Area \Delta = \sqrt{s(s-a)(s-b)(s-c)}\] where s is the semi perimeter of the triangle: \[s = \frac{ a +b +c }{ 2}\] This method is called heron's method 

OpenStudy (anonymous):

ANd is base and height is known the diagram will be: 

OpenStudy (anonymous):

@lovetodowork your question's answered

OpenStudy (juanpablojr):

A=\[\frac{ 1 }{ 2}\times b \times h\]